# Explicit uniform lower bounds for the canonical height on elliptic curves over abelian extensions

### v1

@@ 1. Introduction
the Néron–Tate height on $\mathcal{E}(\bar{F})$ (see [SilvermanArithmeticEC, Definition on p. 248] for more details). We write $\mathcal{E}_{tors}$ for the torsion subgroup of $\mathcal{E}$, and for an integer $n\geq 1$, we denote by $\mathcal{E}[n]$ the subgroup of $n$-torsion points.
@@ 2. Auxiliary results
(2) $\hat{h}(P)\geq\frac{1}{2(a^{2}+b^{2})}\left(\frac{\log p\cdot[F_{\wp}:\mathbb{Q}_{p}]}{\rho[F:\mathbb{Q}]}-C_{1}\right),$
@@
$C_{1}=\frac{1}{6}\left(\log 2+\frac{22}{3}+\log\left(\max\left(1,|j_{\mathcal{E}}|\right)\right)\right),$
@@
$\sum_{v\in M_{L}^{\infty}}d_{v}\lambda_{v}(aP-b\sigma(P))$ $=\sum_{v\in M_{\mathbb{Q}}^{\infty}}\sum_{w\mid v}d_{w}\lambda_{w}(aP-b\sigma(P))$
$\geq-C_{1}\sum_{v\in M_{\mathbb{Q}}^{\infty}}\sum_{w\mid v}d_{w}$
$=-C_{1}$
where the last equality is well known (see for instance [bombieri2006heights, Corollary 1.3.2]) and $C_{1}$ is the constant in the statement.
@@
$\hat{h}(aP-b\sigma(P))\geq\frac{\log p\cdot[F_{\wp}:\mathbb{Q}_{p}]}{\rho[F:\mathbb{Q}]}-C_{1}.$
@@
$2(a^{2}+b^{2})\hat{h}(P)\geq\frac{\log p\cdot[F_{\wp}:\mathbb{Q}_{p}]}{\rho[F:\mathbb{Q}]}-C_{1}.$
@@
$\hat{h}(P)\geq\frac{1}{4q^{4}}\left(\frac{1}{d}\log q-C_{1}\right),$
where $d=[F:\mathbb{Q}]$ and
$C_{1}=\frac{1}{6}\left(\log 2+\frac{22}{3}+\log\left(\max\left(1,|j_{\mathcal{E}}|\right)\right)\right),$
with $j_{\mathcal{E}}$ being the $j$-invariant of $\mathcal{E}$.
@@
Let $\mathfrak{Q}$ be a prime ideal of $L$ over $\wp$. We define the decomposition group and ramification groups as follows:
@@
Suppose that there exists $\tau\in G_{0}$ such that $qP\neq q\tau(P)$. Since $\tau(P)-P\in\mathcal{E}(\mathfrak{Q})$, and $e\leq q$, we deduce that
$q(\tau(P)-P)\in\mathcal{E}(\mathfrak{Q}^{q})\subseteq\mathcal{E}(\mathfrak{Q}^{e}).$
@@
$\hat{h}(P)\geq\frac{1}{4q^{2}}\left(\frac{\log p\cdot[F_{v}:\mathbb{Q}_{v}]}{e(\wp/p)[F:\mathbb{Q}]}-C_{1}\right).$
@@
$\hat{h}(P)\geq\frac{1}{4q^{2}}\left(\frac{1}{d}\log q-C_{1}\right).$
@@
$\hat{h}(P)\geq\frac{1}{2q^{2}(q^{2}+1)}\left(\frac{1}{d}\log q-C_{1}\right).$
@@
$\hat{h}(P)\geq\frac{1}{4q^{2}}\left(\frac{1}{d}\log q-C_{1}\right).$
@@
Set $T:=\tau(Q)-Q$ and $R:=\tau(P)-P$. Let $K$ be the complex multiplication field of the elliptic curve $\mathcal{E}$. Let’s assume for the moment that $F$ contains the Hilbert class field of $K$.
@@
$\hat{h}(Q)\geq\frac{1}{4q^{4}}\left(\frac{1}{d}\log q-C_{1}\right),$
@@
$\hat{h}(P)$ $\geq\left(8(4+(g_{2}-g_{1})^{2})q^{4}\right)^{-1}\left(\frac{1}{d}\log q-C_{1}\right)$
$\geq\left((5+27d+162d^{2})4^{3}q^{4}\right)^{-1}\left(\frac{1}{d}\log q-C_{1}\right).$
@@
$2e^{dC_{1}}\leq p\leq q\leq p^{d}<4^{d}e^{d^{2}C_{1}}.$
Therefore,
$\left((5+27d+162d^{2})4^{3}q^{4}\right)^{-1}\left(\frac{1}{d}\log q-C_{1}\right)\geq\left((5d+27d^{2}+162d^{3})4^{4d+4}e^{4d^{2}C_{1}}\right)^{-1}.$
Recalling that $C_{1}=\frac{1}{6}\left(\log 2+\frac{22}{3}+\log\left(\max\left(1,|j_{\mathcal{E}}|\right)\right)\right)$, and observing that $5d+27d^{2}+162d^{3}\leq 2^{8d}$, we finally have
$\hat{h}(P)$ $\geq\left((5d+27d^{2}+162d^{3})2^{8d+8}e^{4d^{2}C_{1}}\right)^{-1}$
$\geq 2^{-6d^{2}C_{1}-16d-8}$
$\geq 2^{-9d^{2}-d^{2}\log(\max(1,|j_{\mathcal{E}}|))-16d-8}$
$\geq 2^{-9d^{2}-16d-8}\max(1,|j_{\mathcal{E}}|)^{-d^{2}}.$
@@ Acknowledgment
The author would like to thank Samuel Le Fourn and Davide Lombardo for providing the proof of Proposition 2.3 and for valuable discussions.

### v2

@@ 1. Introduction
the Néron–Tate height on $\mathcal{E}(\bar{F})$ (see [SilvermanArithmeticEC, Definition on p. 248] for more details). We write $\mathcal{E}_{tors}$ for the torsion subgroup of $\mathcal{E}$, and for an integer $n\geq 1$, we denote by $\mathcal{E}[n]$ the subgroup of $n$-torsion points. For an algebraic number $\alpha$, we denote by $h(\alpha)$ its absolute logarithmic Weil height (see [bombieri2006heights, Definition 1.5.4] for more details)
@@ 2. Auxiliary results
(2) $\hat{h}(P)\geq\frac{1}{2(a^{2}+b^{2})}\left(\frac{\log p\cdot[F_{\wp}:\mathbb{Q}_{p}]}{\rho[F:\mathbb{Q}]}-C(j_{\mathcal{E}})\right),$
@@
$C(j_{\mathcal{E}})=\frac{1}{6}\left(\log 2+\frac{22}{3}+h\left(j_{\mathcal{E}}\right)\right),$
@@
$\sum_{v\in M_{L}^{\infty}}d_{v}\lambda_{v}(aP-b\sigma(P))$ $\geq-\sum_{v\in M_{L}^{\infty}}\frac{d_{w}}{6}\left(\log 2+\frac{22}{3}+\log^{+}|j_{\mathcal{E}}|_{w}\right)$
where $\log^{+}|j_{\mathcal{E}}|_{w}=\log\left(\max\left(1,|j_{\mathcal{E}}|_{w}\right)\right)$. Following [bombieri2006heights, Corollary 1.3.2] we have $\sum_{v\in M_{L}^{\infty}}d_{w}=1,$ and since $j_{\mathcal{E}}$ is an algebraic integer (see [AdvancedTopicEllipticCurve, Theorem 6.1]), we also have
$\sum_{v\in M_{L}^{\infty}}d_{w}\log^{+}|j_{\mathcal{E}}|_{w}=h(j_{\mathcal{E}}).$
Thus, we have
$\sum_{v\in M_{L}^{\infty}}d_{v}\lambda_{v}(aP-b\sigma(P))\geq-\frac{1}{6}\left(\log 2+\frac{22}{3}\right)-\frac{1}{6}h(j_{\mathcal{E}})=-C(j_{\mathcal{E}}).$
@@
$\hat{h}(aP-b\sigma(P))\geq\frac{\log p\cdot[F_{\wp}:\mathbb{Q}_{p}]}{\rho[F:\mathbb{Q}]}-C(j_{\mathcal{E}}).$
@@
$2(a^{2}+b^{2})\hat{h}(P)\geq\frac{\log p\cdot[F_{\wp}:\mathbb{Q}_{p}]}{\rho[F:\mathbb{Q}]}-C(j_{\mathcal{E}}).$
@@
$\hat{h}(P)\geq\frac{1}{4q^{4}}\left(\frac{1}{d}\log q-C(j_{\mathcal{E}})\right),$
where $d=[F:\mathbb{Q}]$ and $C(j_{\mathcal{E}})$ is as in Lemma 2.1, with $j_{\mathcal{E}}$ being the $j$-invariant of $\mathcal{E}$.
@@
Let $\mathcal{O}_{L}$ be the ring of integers of $L$. Let $\mathfrak{Q}$ be a prime ideal of $\mathcal{O}_{L}$ over $\wp$. We define the decomposition group and ramification groups as follows:
@@
Suppose that there exists $\tau\in G_{0}$ such that $qP\neq q\tau(P)$. Since $\tau(P)-P\in\mathcal{E}(\mathfrak{Q})$, meaning that $\tau(P)-P$ lies in the kernel of the reduction map $E(\mathcal{O}_{L})\longrightarrow E(\mathcal{O}_{L}/\mathfrak{Q}),$ in the sense of [SilvermanArithmeticEC, Chap. VII.2], and since $e\leq q$, we obtain
$q\left(\tau(P)-P\right)\in\mathcal{E}(\mathfrak{Q}^{q})\subseteq\mathcal{E}(\mathfrak{Q}^{e}).$
@@
$\hat{h}(P)\geq\frac{1}{4q^{2}}\left(\frac{\log p\cdot[F_{v}:\mathbb{Q}_{v}]}{e(\wp/p)[F:\mathbb{Q}]}-C(j_{\mathcal{E}})\right).$
@@
$\hat{h}(P)\geq\frac{1}{4q^{2}}\left(\frac{1}{d}\log q-C(j_{\mathcal{E}})\right).$
@@
$\hat{h}(P)\geq\frac{1}{2q^{2}(q^{2}+1)}\left(\frac{1}{d}\log q-C(j_{\mathcal{E}})\right).$
@@
$\hat{h}(P)\geq\frac{1}{4q^{2}}\left(\frac{1}{d}\log q-C(j_{\mathcal{E}})\right).$
@@
Set $T:=\tau(Q)-Q$ and $R:=\tau(P)-P$. Let $K$ be the complex multiplication field of the elliptic curve $\mathcal{E}$. Let’s assume for the moment that $F$ contains the Hilbert class field of $K$. This ensures by [AdvancedTopicEllipticCurve, Theorem II.2.3] that the group $\mathrm{Gal}(L(E[q^{\prime}])/F)$ is abelian.
@@
$\hat{h}(Q)\geq\frac{1}{4q^{4}}\left(\frac{1}{d}\log q-C(j_{\mathcal{E}})\right),$
@@
$\hat{h}(P)$ $\geq\left(8(4+(g_{2}-g_{1})^{2})q^{4}\right)^{-1}\left(\frac{1}{d}\log q-C(j_{\mathcal{E}})\right)$
$\geq\left((5+27d+162d^{2})4^{3}q^{4}\right)^{-1}\left(\frac{1}{d}\log q-C(j_{\mathcal{E}})\right).$
@@
$2e^{dC(j_{E})}\leq p\leq q\leq p^{d}<4^{d}e^{d^{2}C(j_{E})}$
and thus
$\tfrac{1}{d}\log q-C(j_{E})\geq\tfrac{1}{d}\log 2\geq\tfrac{1}{4d}.$
Therefore, taking into account $(5+27d+162d^{2})d\leq 2^{8d}$ and $e^{4}\leq 2^{6}$,
$\hat{h}(P)$ $\geq\big((5+27d^{2}+162d^{3})4^{3}q^{4}\big)^{-1}\big(\tfrac{1}{d}\log q-C(j_{E})\big)$
$\geq\big((5+27d^{2}+162d^{3})d4^{4d+4}e^{4d^{2}C(j_{E})}\big)^{-1}$
$\geq 2^{-16d-8-6d^{2}C(j_{E})}.$
By definition of $C(j_{E})$ we have
$2^{6C(j_{E})}=2^{\log 2+22/3+h(j_{E})}\leq 2^{9+h(j_{E})}.$
Thus we finally get:
$\hat{h}(P)\geq 2^{-(9+h(j_{E}))d^{2}-16d-8}.$
@@ Acknowledgment
The author would like to thank Samuel Le Fourn and Davide Lombardo for providing the proof of Proposition 2.3 and for valuable discussions, and also Gabriel Dill for pointing out several inaccuracies.